TWO RESISTANCES , R1=(16+_0.3) AND R2 = (48+_0.5 ) ARE CONNECTED in parallel. FIND THE 1)TOTAL RESISTANCE OF THE COMBINATION 2)MAXIMUM PERCENTAGE ERROR

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
Find the first Nominal values of Resistances simply .i.i.e for 16 and 48 ohm then calculate error separately.
in series Rnet = R1 + R2 =  16 + 48 = 64 ohm
for error take partial differnetiaion both sides.
d(Rnet) =d(R1) + d(R2) 
            = 0.3 + 0.5 = 0.8
then net resistance with eroor is 64+0.8 =64.8
we always take maximum error thus with plus sign.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)